A 21 nF capacitor is connected across an AC generator that produces a peak voltage of 4.0 V.

At what frequency f is the peak current 44 mA?
Answer #1

Vc = ( 1/(2\pif*C) ) * Im

Vc/Im = ( 1/(2\pif*C) )

Im/Vc = 2\pif*C

f = (Im/Vc) (1/(2\piC))

f = ( 44*10-3 A / 4.0 V) ( 1 / ( 2\pi*21*10-9 F )
f = 8.33*104
Hz